OpenStudy (anonymous):

Solve for the variable. p - 1 = - 54 p =

OpenStudy (anonymous):

p=-53

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

Solve for the variable. d - (-6) = 15 d =

OpenStudy (texaschic101):

d - (-6) = 15 d + 6 = 15 d = 15 - 6 d = 9

OpenStudy (anonymous):

Choose the operation that will isolate the variable in this expression. 3p = -27 Multiply both sides by 3. Divide both sides by 3. Add 27 to both sides.

OpenStudy (texaschic101):

divide both sides by 3